Husky Enterprises recently sold an issue of 13-year maturity bonds. The bonds were sold at a deep discount price of $375 each. After flotation costs, Husky received $358.41 each. The bonds have a $1,000 maturity value and pay $35 interest at the end of each year. Compute the after-tax cost of debt for these bonds if Husky’s marginal tax rate is 40 percent. Round your answer to two decimal places. 10.12% would be entered as 10.12
Answer #1

Using financial calculator
Input: FV= 1000

PMT = 35

PV = price after flotation = -358.41

N = 13

Solve for I/Y as 14.98

After tax cost = YTM*(1-Tax)

= 14.98*(1-0.4)

=8.99
Answer #2

To calculate the after-tax cost of debt for the bonds, we need to consider both the interest payment and the flotation costs.

  1. Interest Payment: The annual interest payment on each bond is $35.

  2. Flotation Costs: The bonds were sold at a deep discount price of $375 each, but Husky received $358.41 after flotation costs per bond. The flotation costs represent the expenses incurred by Husky to issue and sell the bonds.

Now, let's calculate the net proceeds per bond (proceeds after flotation costs) as follows:

Net Proceeds per Bond = Selling Price - Flotation Costs Net Proceeds per Bond = $375 - $358.41 = $16.59

Next, we need to calculate the after-tax interest expense on the bonds. Since interest expense is tax-deductible, the after-tax interest expense is reduced by the tax savings. The tax savings is equal to the interest expense multiplied by the marginal tax rate (40% in this case).

After-Tax Interest Expense = Interest Expense - (Interest Expense * Marginal Tax Rate) After-Tax Interest Expense = $35 - ($35 * 0.40) = $35 - $14 = $21

Now, we can calculate the after-tax cost of debt using the following formula:

After-Tax Cost of Debt = (After-Tax Interest Expense + Flotation Costs) / Net Proceeds per Bond

After-Tax Cost of Debt = ($21 + $358.41) / $16.59 ≈ $379.41 / $16.59 ≈ 22.86

Finally, convert the result to a percentage:

After-Tax Cost of Debt = 22.86%

So, the after-tax cost of debt for these bonds is approximately 22.86%.

answered by: Hydra Master